If 37 grams of methanol is produced from the reaction of 60 grams of CO and 6 grams of H2.Determine the reagent,excess reagent, percent yield and theoretical yield of reaction


1
Expert's answer
2022-04-14T03:03:03-0400

CO + 2H2 = CH3OH

n(H2)=m(H2)/Mr(H2)=6g / 4g/moll = 1.5 moles

n(CO) = m(CO)/Mr(CO) = 60g / 28g/moll = 2.14 moles

CO excess reagent

Hlimiting reagent


m(CH3OH) = ( m(H2)*Mr(CH3OH) ) / Mr(H2) = (6g * 32g/moll) / 4g/moll = 48g - theoretical yield


percent yield = (37/48)*100% = 77.1%